OpenStudy (anonymous):

What is the geometric mean of 3 and 15?

hero (hero):

In general the geometric mean of a number is \(ab = x^2\) Sometimes it is written like this: \(\large\frac{a}{x} = \frac{x}{b} \) In this case, a = 3 , b = 5 so \( 3 \dot\ 5 = x^2\)

hero (hero):

To find the geometric mean, solve for x

OpenStudy (anonymous):

THANKS..